How many atoms are in
117
g
of calcium?
1
Expert's answer
2020-03-31T12:41:37-0400

Hello



N(Ca)=n(Ca)*NA

NA=6.02*1023


n(Ca)=m(Ca)/M(Ca)=117/40=2.925mole

N(Ca)=2.925*6.02*1023=17.609*1023atoms


Answer - 17.609*1023atoms